OpenStudy (anonymous):

The point given is on the terminal side of an angle theta. Find the exact values of the six trig functions of theta. (-3,2)

OpenStudy (anonymous):

sin = 2/sqrt13 cos = -3/sqrt13 tan = -2/3 csc = sqrt13/2 sec = -sqrt13/3 cot = -3/2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

sorries i dont know how to do this im just looking for the explanation :P

OpenStudy (anonymous):

No worries...you take the point and find r

OpenStudy (anonymous):

r = sqrt x^2+y^2

OpenStudy (anonymous):

so... r=sqrt -3^2 + 2^2

OpenStudy (anonymous):

r=sqrt13

OpenStudy (anonymous):

then you just plug x,y,and r in into the six trig functions of theta accordingly.

OpenStudy (anonymous):

i dont like trig that much thats why :P

OpenStudy (anonymous):

but thanks :)